Which value is the smallest?

A. 4.5 x 102 mm
B. 450 cm
C. 4.5 cm
D. 4.5 m
E. 4.5 x 106 nm
Answer #1

First, Let us convert all the unit into a single unit,meter

Some Conversion to remember:

1 mm = 10-3 m

1 cm = 10-2 m

1 nm (nanometer) = 10-9 m

Now let us go by each option:

(A).   4.5 x 102 mm = 4.5 x 102 x10-3 m =  4.5 x 10-1m= 0.45m

(B). 450 cm = 450x 10-2m = 4.5 m

(C). 4.5 cm = 4.5x 10-2m = 0.45m

(D). 4.5m (given)

(E). 4.5 x 106 nm = 4.5 x 106x 10-9 m = 4.5 x 10-3m = 0.0045m

So, we can see that 4.5x 106nm = 0.0045m , is the smallest value.

Please give it a thumbsup if you like my answer. And feel free to ask in comment box if you have any doubt regarding solution